OpenStudy (anonymous):

Write the equation for the inverse of the function y = pi/4 + sin x.

OpenStudy (anonymous):

Isolate x

OpenStudy (ranga):

To find the inverse function, replace x with y and y with x and then solve for y.

OpenStudy (anonymous):

I've got that far, I just don't know how to solve for y once I get there. I got x = arcsin (y - pi/4)

OpenStudy (ranga):

y = pi/4 + sin x To find the inverse function, First step: replace x with y and y with x x = pi/4 + sin(y) solve for y. subtract pi/4 from both sides x - pi/4 = sin(y) take arcsin on both sides arcsin(x-pi/4) = y y = arcsin(x-pi/4) is the inverse function.

OpenStudy (ranga):

Or in your answer you can replace x with y and y with x and you will get the same answer. You are doing the interchange as a last step. I prefer interchanging x and y as the first step in finding the inverse.

OpenStudy (anonymous):

thank you guys for the help

OpenStudy (ranga):

you are welcome.